4. Find the orthogonal decomposition of v with respect to W A. --[ : ) ----([1][:]B. 0 0 l W span ( 5 )
Answer #1

4] A. 4 -2 V wespan -([:] [:] 3 Let V - V2 = - (+] V, . r = (1) + 2 (7) +(!) (1) tra id) { ww his an orthogonal basis of w --([3) - (:) . [ پد = 솔 لا - % - اد : 441 - .: the oxthogonal de composition of v with respect to wis و [4]3 B] VE W-span Let VI- [ V2 = D 1 V, V2= (1) (1) + )(0) + (11-1) + 0) (1) = 1 -10 Ev, TV₂ {v, V, } is an PROI) = VRV orthogon